1.9) Circle the oxidation state of the carbon atom in the molecule dichloromethane (CH2Cl2)? a) -2 b)-1 c) 0 d) +1 e) +2
media%2F727%2F727e4b59-0b96-4a0d-9d9b-f2
Answer #1

1.9)
H is attached to C
so, oxidation state of H is +1

oxidation state of Cl is -1

let oxidation state of C be x
use,
x + 2*(oxidation state of H) + 2*(oxidation state of Cl) = 0
x + 2*1 + 2*(-1) = 0
x + 2 - 2 = 0
x = 0
so,
oxidation state of C = 0

Answer : option c

i am allowed to do only one question at a time